# Scalable In-Context Q-Learning

### v1

@@ 5.5 Visualization Insights
We gain deep insights into the prompt construction process through t-SNE visualization on Cheetah-Vel and Ant-Dir tasks, as shown in Fig. 5. We use a continuous color spectrum to indicate task similarity. The initially entangled transitions are transformed into well-separated clusters in the prompt space, where points from different tasks are more clearly distinguished and similar tasks are grouped more closely. Further, prompt representations of Cheetah-Vel form a clear rectilinear distribution from blue (low velocity) to red (high velocity), exactly aligning with the rectilinear spectrum of target velocities in a physical sense. Similarly, prompt representations of Ant-Dir follow a cyclic spectrum that matches the periodicity of angular directions in a physical sense. This interesting finding highlights SICQL’s ability to distill meaningful task-specific information from raw transitions, enabling precise task inference to facilitate in-context learning capacity.
@@
Figure 5: t-SNE visualization on Cheetah-Vel and Ant-Dir, where tasks differ in target velocities of $[0.1,3.0]$ and target directions of $[0,2\pi]$. Data representations of raw transitions $(s,a,r,s^{\prime})$ and precise prompts $\beta$ from a distribution of tasks are mapped into rainbow-colored points.
@@ 6 Conclusions, Limitations, and Future Work
In the paper, we propose SICQL, an innovative framework that introduces dynamic programming and world modeling to enable fundamental reward maximization and efficient task generalization in ICRL. SICQL employs a multi-head transformer to jointly predict optimal policies and in-context value functions, guided by a pretrained world model that encodes precise task-relevant information for efficient prompt construction. Policy improvement is achieved by fitting in-context value functions with expectile regression and distilling them into policy extraction via advantage-weighted regression, allowing for reward maximization while retaining the scalability and stability of supervised pretraining. Extensive evaluations verify the consistent superiority of SICQL over a range of competitive baselines.
Though, the length of our prompt is the same as the sampled transitions, which may not be short enough, especially for long-horizon interactive problems. Future work can explore extracting task information with more compact structures, e.g., encoding an episode or a skill into a single token. Another promising step is to leverage natural language as a higher-level task prompt for ICLR, harnessing a wealth of knowledge within pretrained language models.
@@ Appendix A Limitations
Leveraging Natural Language for SICQL: A promising direction is using natural language as a higher-level prompt for SICQL. Pretrained language models could enhance knowledge transfer, improving generalization and adaptability, offering valuable opportunities for future research.
@@ Appendix B Algorithm Pseudocodes
Based on the implementations presented in Sec. 4, this section provides an overview of the procedural steps of our method. Initially, Algorithm 1 outlines the pretraining process for the world model. Subsequently, Algorithm 2 describes the pipeline for training and testing SICQL.
@@
       Compute the predicted reward and next state$[\hat{r}_{t},\hat{s}_{t+1}]\!=\!D_{\varphi}(s_{t},a_{t};z_{t}^{i})$ Update $E_{\phi}$ and $D_{\varphi}$ using the loss as $\mathcal{L}(\phi,\varphi)=\mathbb{E}_{\eta_{t}^{i}\sim M_{i}}\!\left[\mathbb{E}_{z^{i}_{t}\sim E_{\phi}(\eta_{t}^{i})}\!\left[\lVert[r_{t},s_{t+1}]-D_{\varphi}(s_{t},a_{t};z_{t}^{i})\rVert^{2}\right]\right]$
@@
      Sample $\left(s_{t}^{i},a_{t}^{i},s_{t+1}^{i},r_{t}^{i}\right)$ and a $h$-step trajectory $[s_{1},a_{1},r_{1},...,s_{h},a_{h},r_{h}]^{i}$ randomly from dataset from $\mathcal{D}^{i}$.
@@ Appendix C The Details of Environments
DarkRoom: The agent is randomly placed in a 10 ×10 grid room, and the goal occupies one of the 100 grid cells. Thus, there are 100 possible goals. The agent’s observation is its current grid cell, i.e., $\mathcal{S}=[10]\times[10]$. At each step, the agent selects one of five actions: move up, down, left, right, or remain stationary. The agent receives a reward of 1 only upon reaching the goal and 0 otherwise. The episode horizon for Dark Room is 100. Consistent with [10], we use 80 of the 100 goals for pretraining and reserve the remaining 20 goals to test our model’s in-context RL capability on unseen tasks.
@@
  • Cheetah-Vel: A planar cheetah must run forward at a specified target velocity along the positive x-axis. Each task is defined by a distinct reward function that penalizes the absolute deviation between the cheetah’s instantaneous velocity and its goal velocity. Goal velocities are drawn uniformly from $[0.1,3.0]$, yielding a suite of tasks with varying targets.
@@
  • Ant-Dir: A quadrupedal ant robot must move in a specified direction. Each task defines a distinct target angle, and the reward is given by the cosine similarity between the agent’s velocity vector and the unit vector in the target direction. Target directions are drawn uniformly from $[0,2\pi]$.
@@
Meta-World: The Meta-Learning 1 (ML1) suite in Meta-World is a minimalist, single-task benchmark for few-shot meta-reinforcement learning. For each task, a Sawyer robotic arm in MuJoCo must reach a randomly Selected target—whose coordinates are withheld from observations—forcing the agent to infer the goal by trial and error. ML1’s sparse information and clear generalization challenge make it one of Meta-World’s most popular robotic manipulation testbeds for evaluating rapid adaptation in meta-RL.
@@
  • Push: A Sawyer arm must push a block to a randomly placed target on a tabletop, with each task varying the block’s start position and corresponding reward function. Agents receive only sparse distance-based feedback and must infer the goal through interaction. This setup evaluates the agent’s ability to explore and adapt its pushing strategy under sparse supervision.
@@ Appendix D The Details of Dataset Construction
Pretraining Datasets for Darkroom: In the Dark Room environment, we gather 1,000 trajectories per task with horizon $H=100$. At each step, we follow the optimal policy with probability $\epsilon$ and a random policy with probability $1-\epsilon$. We choose $\epsilon$ so that the mean return of the pretraining datasets is below 40% of the optimal policy return, reflecting the challenging yet common scenarios.
Pretraining Datasets for MuJoCo and Meta-World: For each evaluation domain, we choose 45 tasks to construct the training datasets and train a single-task policy independently for each task. We use soft actor-critic (SAC) [40] for the MuJoCo domains. We use Proximal Policy Optimization (PPO) [41] for the Meta-World domain. We collect two types of offline datasets for each evaluation domain as
@@ Appendix E The Details of Baselines
This section presents seven representative baselines addressing the meta-task generalization problem, including one context-based offline meta-reinforcement learning (COMRL) method and six ICRL approaches. These baselines are thoughtfully selected to span the major domains of current offline meta-task research. Furthermore, since our proposed SICQL method belongs to the ICRL category, we incorporate more methods from this class as baselines for a comprehensive comparison. The detailed descriptions of these baselines are as follows:
@@
  • IC-IQL [16] extends standard ICRL by explicitly optimizing reinforcement learning objectives instead of relying solely on supervised losses as in AD. It augments AD with a Q-learning loss, where the value function is trained by updating both its head and the shared transformer backbone, while the policy is trained by updating only its own head.
@@
  • DIT [14] addresses the limitations of standard autoregressive imitation learning when trained on suboptimal trajectories. Instead of treating trajectory prediction purely as supervised learning, DIT emulates an actor-critic algorithm in-context by applying a weighted maximum likelihood estimation (WMLE) loss, where the weights are directly computed from observed rewards.
@@
  • DICP  [12], combines model-based reinforcement learning with prior in-context RL approaches such as AD and IDT. DICP leverages a pre-trained Transformer not only to condition past transitions but also to simulate future trajectories and estimate long-term returns, enabling proactive decision-making without parameter updates. DICP jointly models environment dynamics and policy improvements in context, resulting in greater adaptability and sample efficiency, particularly in long-horizon tasks.
@@
  • IDT  [13], employs a hierarchical learning framework that decomposes decision-making across temporal scales to address the high computational cost of prior in-context RL methods on long-horizon tasks. Its architecture consists of three modules: (i) Decision-Making, which predicts high-level decisions; (ii) Decision-to-Go, which decodes these decisions into low-level actions; and (iii) Decision-Review, which maps low-level actions back to high-level representations. Built upon a transformer architecture similar to that of AD, IDT addresses complex decision-making in long-horizon tasks while reducing computational costs through hierarchical modeling.
@@
  • AD  [9], casts in-context RL as a supervised sequence-modeling task: a causal Transformer is pretrained on across-episodic trajectories covering the entire RL learning to predict subsequent actions, thereby emulating standard RL update dynamics without explicit gradient updates. In AD, trajectories gathered across episodes are organized into fixed-length sequences of length $H$, with each trajectory tokenized as an interleaved series of states, actions, and rewards. This allows the model to learn purely from contextual information, enabling in-context adaptation of policies to new tasks.
@@
  • DPT  [10], adopts a supervised training paradigm to enable in-context learning for RL tasks. The core idea is to train a transformer to predict the provided optimal action for a given query state purely based on context, using interaction histories from diverse tasks as in-context datasets. DPT treats each transition tuple $(s,a,s^{\prime},r)$ as a single token, rather than decomposing it into separate embeddings. This allows the attention mechanism to directly model relationships between full transitions, preserving the structural and semantic integrity of the interaction data.
@@
  • UNICORN [42], integrates representative methods such as FOCAL [43], CORRO [22], and CSRO [44], this work proposes a unified information-theoretic framework that interprets these approaches as optimizing different approximation bounds of the mutual information between task variables and their latent representations. Building on the information bottleneck principle, it further derives a general and unified objective for task representation learning, facilitating the extraction of robust and transferable task representations. This method provides a unified information-theoretic perspective that summarizes and connects several mainstream context-based offline meta-RL approaches. As it integrates key ideas from representative COMRL algorithms, it can serve as a strong and generalizable baseline.
@@
Mainstream ICRL algorithms predominantly rely on imitation learning paradigms, which tend to learn suboptimal behaviors when pretraining data is limited in size or quality. To ensure fair comparison, we standardize the dataset size and quality across methods during pretraining. In our task setting, AD struggles to extract effective policy improvement operators due to insufficient offline data. Therefore, we adopt an AD variant that incorporates a reward-based trajectory sorting mechanism from AT [45] to better distill policy improvements. Additionally, since UNICORN under the COMRL setting requires warm-up data for task representation inference, we align it with the ICRL setup by initializing task representations using randomly sampled trajectories and updating them after each episode.
Appendix F Implementation Details of SICQL
@@
World Model. In this paper, we adopt streamlined architectures for each component of the world model: the context encoder, and the dynamics decoder. The context encoder first employs a fully connected multi-layer perceptron (MLP) followed by a gated recurrent unit (GRU) network, both using ReLU activations. The GRU processes the agent’s $k$ step history $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$. and produces a 128-dimensional hidden vector. This vector is then projected via the MLP into a 16-dimensional task representation $z_{t}^{i}$. The reward decoder is an MLP that receives the tuple $(s_{t}^{i},a_{t}^{i},s_{t+1}^{i},z_{t}^{i})$ and passes it through two hidden layers of size 128 to predict the scalar reward $\hat{r}_{t}^{i}$. Analogously, the state transition decoder is an MLP that takes $(s_{t}^{i},a_{t}^{i},z_{t}^{i})$ as input and uses two 128-dimensional hidden layers to predict the next state $\hat{s}_{t+1}^{i}$.
[…]
Table 8: The network configurations used for SICQL.
@@
Algorithm Hyperparameters. We evaluate the proposed SICQL algorithm on six environments: Darkroom, Push, Reach, Cheetah-Vel, Walker-Param, and Ant-Dir. For all experiments, we use the Adam optimizer with a weight decay of 1e-4, gradient-norm clipping at 10, an experience step of 4, and a batch size of 128. Table 9 summarizes the detailed hyperparameters of SICQL in each domain.
Table 9: Hyperparameters of SICQL on various domains.
@@
Appendix G Graphic Results for Hyperparameter Analysis
Due to limited space, we only present tabular results in Sec. 5.4 in the main text. Here, we present the corresponding graphic results. Figure 6 shows the few-shot evaluation return curves of SICQL on Mixed datasets with varying values of $\lambda$. Figure 7 presents few-shot evaluation return curves of SICQL on Mixed datasets with varying values of $k$.
Figure 6: Few-shot evaluation return curves of SICQL on Mixed datasets with varying values of $\lambda$.
Figure 7: Few-shot evaluation return curves of SICQL on Mixed datasets with varying values of $k$.

### v2

@@ 4.6 Visualization Insights
We gain deep insights into the prompt construction process through t-SNE visualization on Cheetah-Vel and Ant-Dir tasks, as shown in Figure 6. We use a continuous color spectrum to indicate task similarity. For each task, we randomly sample 200 transitions and encode them into precise task representations using the pretrained world model. The initially entangled transitions are transformed into well-separated clusters in the prompt space, where points from different tasks are clearly distinguished and similar tasks are grouped more closely. Further, prompt representations of Cheetah-Vel form a clear rectilinear distribution from blue (low velocity) to red (high velocity), exactly aligning with the rectilinear spectrum of target velocities in a physical sense. Similarly, prompt representations of Ant-Dir follow a cyclic spectrum that matches the periodicity of angular directions in a physical sense. This interesting finding highlights S-ICQL’s ability to harness world modeling for distilling meaningful task-specific information from raw transitions, enabling precise task inference to facilitate in-context learning capacity.
@@
Figure 6: t-SNE visualization on Cheetah-Vel and Ant-Dir, where tasks differ in target velocities of $[0.1,3.0]$ and target directions of $0,2\pi]$. Data representations of raw transitions $(s,a,r,s^{\prime})$ and precise prompts $\beta$ from a distribution of tasks are mapped into rainbow-colored points.
@@ 5 Conclusions, Limitations, and Future Work
In the paper, we propose S-ICQL, an innovative framework that introduces dynamic programming and world modeling to enable fundamental reward maximization and efficient task generalization in ICRL. S-ICQL employs a multi-head transformer to jointly predict optimal policies and in-context value functions, guided by a pretrained world model that encodes precise task-relevant information for efficient prompt construction. Policy improvement is achieved by fitting in-context value functions with expectile regression and extracting policies via advantage-weighted regression, enabling reward maximization while preserving the scalability and stability of supervised pretraining. Extensive evaluations verify the consistent superiority of S-ICQL over a range of competitive baselines.
Though, the length of our prompt is the same as the sampled transitions, which may not be short enough, especially for long-horizon interactive problems. Future work can explore extracting task information with more compact structures, e.g., encoding an episode or a skill into a single token. Another promising step is to leverage natural language as a higher-level task prompt for ICLR, harnessing a wealth of knowledge within pretrained large language models (LLMs).
Ethics Statement
We are not aware of any major ethical concerns arising from our work. Our study is conducted entirely within RL benchmarking environments, using only publicly available models and datasets for training and evaluation. No human subjects were involved, and our research does not introduce sensitive or potentially harmful insights.
Reproducibility statement
We provide detailed descriptions of our experimental settings in the Appendix, including environments, datasets, implementation details, and hyperparameters. In addition, we provide the original code in supplementary materials to facilitate reproducibility and verification of our results.
@@ Appendix A Limitations
Leveraging Natural Language for S-ICQL: A promising direction is using natural language as a higher-level prompt for S-ICQL. Pretrained language models could enhance knowledge transfer, improving generalization and adaptability, offering valuable opportunities for future research.
@@ Appendix B Algorithm Pseudocodes
Based on the implementations presented in Sec. 3, this section provides an overview of the procedural steps of our method. Initially, Algorithm 1 outlines the pretraining process for the world model. Subsequently, Algorithm 2 describes the pipeline for training and testing S-ICQL.
@@
[…]
@@
  Sample $\left(s_{t}^{i},a_{t}^{i},s_{t+1}^{i},r_{t}^{i}\right)$ and a $h$-step trajectory $s_{1},a_{1},r_{1},...,s_{h},a_{h},r_{h}]^{i}$ randomly from dataset from $\mathcal{D}^{i}$.
@@ Appendix C The Details of Environments
DarkRoom: The agent is randomly placed in a 10 ×10 grid room, and the goal occupies one of the 100 grid cells. Thus, there are 100 possible goals. The agent’s observation is its current grid cell, i.e., $\mathcal{S}=[10]\times[10]$. At each step, the agent selects one of five actions: move up, down, left, right, or remain stationary. The agent receives a reward of 1 only upon reaching the goal and 0 otherwise. The episode horizon for Dark Room is 100. Consistent with (Lee et al., 2023), we use 80 of the 100 goals for pretraining and reserve the remaining 20 goals to test our model’s in‑context RL capability on unseen tasks.
@@
  • Cheetah-Vel: A planar cheetah must run forward at a specified target velocity along the positive x‑axis. Each task is defined by a distinct reward function that penalizes the absolute deviation between the cheetah’s instantaneous velocity and its goal velocity. Goal velocities are drawn uniformly from $[0.1,3.0]$, yielding a suite of tasks with varying targets.
@@
  • Ant-Dir: A quadrupedal ant robot must move in a specified direction. Each task defines a distinct target angle, and the reward is given by the cosine similarity between the agent’s velocity vector and the unit vector in the target direction. Target directions are drawn uniformly from $0,2\pi]$.
@@
Meta-World: The Meta‑Learning 1 (ML1) suite in Meta‑World is a minimalist, single‑task benchmark for few‑shot meta‑reinforcement learning. For each task, a Sawyer robotic arm in MuJoCo must reach a randomly Selected target—whose coordinates are withheld from observations—forcing the agent to infer the goal by trial and error. ML1’s sparse information and clear generalization challenge make it one of Meta‑World’s most popular robotic manipulation testbeds for evaluating rapid adaptation in meta‑RL.
@@
  • Push: A Sawyer arm must push a block to a randomly placed target on a tabletop, with each task varying the block’s start position and corresponding reward function. Agents receive only sparse distance‑based feedback and must infer the goal through interaction. This setup evaluates the agent’s ability to explore and adapt its pushing strategy under sparse supervision.
@@ Appendix D The Details of Dataset Construction
Pretraining Datasets for Darkroom: In the Dark Room environment, the horizon of each trajectory is set to $100$ steps. At each step, we follow the optimal policy with probability $\epsilon$ and a random policy with probability $1-\epsilon$. We choose $\epsilon$ so that the mean return of the pretraining datasets is below 40% of the optimal policy return, reflecting the challenging yet common scenarios.
Pretraining Datasets for MuJoCo and Meta-World: For each evaluation domain, we choose 45 tasks to construct the training datasets and train a single-task policy independently for each task. We use soft actor-critic (SAC) (Haarnoja et al., 2018) for the MuJoCo domains. We use Proximal Policy Optimization (PPO) (Schulman et al., 2017) for the Meta-World domain. We collect two types of offline datasets for each evaluation domain as
@@ Appendix E The Details of Baselines
This section presents seven representative baselines addressing the meta-task generalization problem, including one context-based offline meta-reinforcement learning (COMRL) method and six ICRL approaches. These baselines are thoughtfully selected to span the major domains of current offline meta-task research. Furthermore, since our proposed S-ICQL method belongs to the ICRL category, we incorporate more methods from this class as baselines for a comprehensive comparison. The detailed descriptions of these baselines are as follows:
@@
  • IC-IQL (Tarasov et al., 2025) extends standard ICRL by explicitly optimizing reinforcement learning objectives instead of relying solely on supervised losses as in AD. It augments AD with a Q-learning loss, both the policy and value are optimized by updating only their heads, with no gradient propagation to the shared transformer backbone.
@@
  • DIT (Dong et al., 2024) addresses the limitations of standard autoregressive imitation learning when trained on suboptimal trajectories. Instead of treating trajectory prediction purely as supervised learning, DIT emulates an actor-critic algorithm in-context by applying a weighted maximum likelihood estimation (WMLE) loss, where the weights are directly computed from observed rewards.
@@
  • DICP  (Son et al., 2025), combines model-based reinforcement learning with prior in-context RL approaches such as AD and IDT. DICP leverages a pre-trained Transformer not only to condition past transitions but also to simulate future trajectories and estimate long-term returns, enabling proactive decision-making without parameter updates. DICP jointly models environment dynamics and policy improvements in context, resulting in greater adaptability and sample efficiency, particularly in long-horizon tasks.
@@
  • IDT  (Huang et al., 2024), employs a hierarchical learning framework that decomposes decision-making across temporal scales to address the high computational cost of prior in-context RL methods on long-horizon tasks. Its architecture consists of three modules: (i) Decision-Making, which predicts high-level decisions; (ii) Decision-to-Go, which decodes these decisions into low-level actions; and (iii) Decision-Review, which maps low-level actions back to high-level representations. Built upon a transformer architecture similar to that of AD, IDT addresses complex decision-making in long-horizon tasks while reducing computational costs through hierarchical modeling.
@@
  • AD  (Laskin et al., 2023), casts in‑context RL as a supervised sequence‑modeling task: a causal Transformer is pretrained on across‑episodic trajectories covering the entire RL learning to predict subsequent actions, thereby emulating standard RL update dynamics without explicit gradient updates. In AD, trajectories gathered across episodes are organized into fixed‐length sequences of length $H$, with each trajectory tokenized as an interleaved series of states, actions, and rewards. This allows the model to learn purely from contextual information, enabling in‑context adaptation of policies to new tasks.
@@
  • DPT  (Lee et al., 2023), adopts a supervised training paradigm to enable in-context learning for RL tasks. The core idea is to train a transformer to predict the provided optimal action for a given query state purely based on context, using interaction histories from diverse tasks as in-context datasets. DPT treats each transition tuple $(s,a,s^{\prime},r)$ as a single token, rather than decomposing it into separate embeddings. This allows the attention mechanism to directly model relationships between full transitions, preserving the structural and semantic integrity of the interaction data.
@@
  • UNICORN (Li et al., 2024b), integrates representative methods such as FOCAL (Li et al., 2021), CORRO (Yuan & Lu, 2022), and CSRO (Gao et al., 2023), this work proposes a unified information-theoretic framework that interprets these approaches as optimizing different approximation bounds of the mutual information between task variables and their latent representations. Building on the information bottleneck principle, it further derives a general and unified objective for task representation learning, facilitating the extraction of robust and transferable task representations. This method provides a unified information-theoretic perspective that summarizes and connects several mainstream context-based offline meta-RL approaches. As it integrates key ideas from representative COMRL algorithms, it can serve as a strong and generalizable baseline.
@@
Mainstream ICRL algorithms predominantly rely on imitation learning paradigms, which tend to learn suboptimal behaviors when pretraining data is limited in size or quality. To ensure fair comparison, we standardize the dataset size and quality across methods during pretraining. In our task setting, AD struggles to extract effective policy improvement operators due to insufficient offline data. Therefore, we adopt an AD variant that incorporates a reward-based trajectory sorting mechanism from AT (Liu & Abbeel, 2023) to better distill policy improvements. Additionally, since UNICORN under the COMRL setting requires warm-up data for task representation inference, we align it with the ICRL setup by initializing task representations using randomly sampled trajectories and updating them after each episode.
Appendix F Implementation Details of S-ICQL
@@
World Model. In this paper, we adopt streamlined architectures for each component of the world model: the context encoder, and the dynamics decoder. The context encoder first employs a fully connected multi‑layer perceptron (MLP) followed by a gated recurrent unit (GRU) network, both using ReLU activations. The GRU processes the agent’s $k$ step history $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$. and produces a 128‑dimensional hidden vector. This vector is then projected via the MLP into a 16‑dimensional task representation $z_{t}^{i}$. The reward decoder is an MLP that receives the tuple $(s_{t}^{i},a_{t}^{i},s_{t+1}^{i},z_{t}^{i})$ and passes it through two hidden layers of size 128 to predict the scalar reward $\hat{r}_{t}^{i}$. Analogously, the state transition decoder is an MLP that takes $(s_{t}^{i},a_{t}^{i},z_{t}^{i})$ as input and uses two 128‑dimensional hidden layers to predict the next state $\hat{s}_{t+1}^{i}$.
Causal Transformer. We implement S-ICQL on top of the official DPT codebase released by DPT (Lee et al., 2023) (https://github.com/jon–lee/decision-pretrained-transformer). We adhere to their architectural design, and construct the embeddings for the GPT-2 backbone as follows. Specifically, given a task dataset $\mathcal{D}^{i}$, we sample $\left(s_{t}^{i},a_{t}^{i},s_{t+1}^{i},r_{t}^{i}\right)$ and a $h$-step trajectory $s_{1},a_{1},r_{1},...,s_{h},a_{h},r_{h}]^{i}$ randomly from dataset from $\mathcal{D}^{i}$. Then we use the pretrained context encoder $E_{\phi}$ to transform raw transitions in this trajectory into precise and lightweight prompt $\beta^{i}:=[z_{1},...,z_{h}]^{i}=E_{\phi}\left([s_{1},a_{1},r_{1},...,s_{h},a_{h},r_{h}]^{i}\right)$. We form state vectors $\xi^{i}_{s_{t}}=(s_{t}^{i},0)$, next state vectors $\xi^{i}_{s_{t+1}}=(s_{t+1}^{i},0)$ and state-action vectors $\xi^{i}_{\{sa\}_{t}}=(s_{t}^{i},a_{t}^{i},0)$ by concatenating the relevant quantities and padding with zeros so that each $\xi^{i}$ has dimension $d_{\xi}:=2d_{S}+d_{A}+1$. The $(h+1)$-length sequence is given by $X=(\xi^{i},z_{1}^{i},\ldots,z_{h}^{i})$. We first apply linear projection $\texttt{Linear}(\cdot)$ to each vector and outputs the sequence $Y=(\hat{y}_{0},\hat{y}_{1},\ldots,\hat{y}_{h})$. Then the $(h+1)$-length tokens are fed into the transformer and predict output autoregressively using a causal self-attention mask. In the output layer, we employ separate linear layers to produce the actions, the state‑values, and the Q‑values, respectively. In summary, Table 7 shows the details of network structures.
Table 7: The network configurations used for S-ICQL.
@@
Algorithm Hyperparameters. We evaluate the proposed S-ICQL algorithm on six environments: Darkroom, Push, Reach, Cheetah‑Vel, Walker‑Param, and Ant‑Dir. For all experiments, we use the Adam optimizer with a weight decay of 1e-4, gradient‑norm clipping at 10, an experience step of 4, and a batch size of 128. Table 8 summarizes the detailed hyperparameters of S-ICQL in each domain.
Table 8: Hyperparameters of S-ICQL on various domains.
@@
Appendix G Hyperparameter Analysis
Table 9: Few-shot converged results of S-ICQL with varying $\lambda$ on Mixed datasets.
$\lambda$ Reach Cheetah-Vel Ant-Dir
[…]
The temperature parameter $\lambda$ in Eq.(6) plays a crucial role in balancing the trade-off between behavior cloning and the greedy pursuit of high Q-values. We conduct experiments to analyze the influence of $\lambda$ on S-ICQL’s performance. Figure 7 and Table 9 present the ablation results across representative domains with varying values of $\lambda$. A smaller $\lambda$ will make the distribution of advantage weights $\exp(\frac{1}{\lambda})(Q-V)$ less uniform, leading to more exploitation of Q-learning. As $\lambda$ decreases from $1.0$, the performance of S-ICQL can be greatly improved, highlighting the essentiality of harnessing Q-learning to steer ICRL architectures toward fundamental reward maximization. Practically, $\lambda=0.01$ or $\lambda=0.001$ leads to a satisfactory performance. The result further confirms the effectiveness and necessity of our in-context Q-learning.
Table 10: Few-shot converged results of S-ICQL with varying $k$ on Mixed datasets.
$K$ Reach Cheetah-Vel Ant-Dir
22 802.00 $\pm$ 10.80 -37.51 $\pm$ 1.80 717.95 $\pm$ 37.81
44 806.97 $\pm$ 6.35 -35.48 $\pm$ 1.33 813.34 $\pm$ 14.12
66 794.19 $\pm$ 7.93 -40.36 $\pm$ 1.36 648.70 $\pm$ 56.75
In prompt construction, the task representation $z_{t}^{i}$ is inferred from the agent’s $k$-step experience $\eta_{t}^{i}$ as $z_{t}^{i}\sim E_{\phi}(\eta_{t}^{i})$. The step $k$ is crucial for capturing task-relevant information. We conduct experiments to investigate the impact of $k$ on S-ICQL’s performance. Figure 8 and Table 10 present the few-shot evaluation returns of S-ICQL across representative environments with varying values of $k$. The results show that S-ICQL’s performance is not sensitive to $k$, with a moderate experience step of $k=4$ performing the best in all evaluated domains. A small $k$ may provide insufficient task-relevant information, limiting precise in-context inference. A large $k$ may introduce redundancy and noise leading to overfitting during world model pretraining. Also, a large $k$ will increase the computation load in pretraining the world model and decrease the speed in constructing the prompt for in-context task inference. In practice, a moderate $k$ achieves fast and precise in-context inference, highlighting the superiority of our prompt design via world modeling.
Figure 7: Few-shot evaluation return curves of S-ICQL on Mixed datasets with varying values of $\lambda$.
Figure 8: Few-shot evaluation return curves of S-ICQL on Mixed datasets with varying values of $k$.
Appendix H Graphic Results for Ablation Study
Due to limited space, we only present tabular results in Sec. 4.2 in the main text. Here, we present the corresponding graphic results. Figure 9 shows the few-shot evaluation return curves of S-ICQL and its ablations on Mixed datasets.
Figure 9: Few-shot evaluation return curves of S-ICQL and its ablations on Mixed datasets. w/o_c removes world modeling, w/o_q removes Q-learning, and w/o_cq removes both components.
Appendix I The Use of Large Language Models (LLMs)
We utilize Large Language Models (LLMs) to assist with polishing the writing and improving text readability. Specifically, LLMs are employed for proofreading, enhancing grammar, and refining sentence structure. The LLM was used solely for editorial purposes to improve clarity and did not contribute to research ideation, experimental design, implementation, analysis, or scientific conclusions. All core research contributions, experiments, and analyses were conducted independently by the authors without LLM assistance.